Determine the point P that partitions AB with A(1,2) B(11,22) into a 1:4 ratio.
1
Expert's answer
2019-02-06T10:37:07-0500

x = 1 + (11-1)/(1+4) = 3

y = 2 + (22-2)/(1+4) = 6

P (3,6)